# The stochastic discrete nonlinear Schrödinger equation: microscopic derivation and finite-temperature phase transition

### v1

@@ Appendix B Phase diagram of the DNSE
The phase diagram of the DNSE, i.e., the bounds for the energy and a representation of the different regimes in the parameter plane, is well established in the literature. The considerations are often based on a standardised version of the DNSE (see, for instance, [rasmussen_statistical_2000, rumpf_simple_2004, BaIuLiVu_PR21]) and those results can be easily translated to our setup using suitable parameter transformations (see e.g. [EBRAHIMI2025134905] for the transformation formulas). For the convenience of the reader, we here summarise the elementary estimates within our notation of eq. (1). The two parameters in our setup are the energy $E=H_{S}$ and the on-site potential strength $\alpha$. While the phase diagram summarises the properties of $H_{S}$ subjected to the constraint $N=1$, the diagram also applies for the SDNSE, eq. (11), if one is able to translate between temperature and energy using the caloric equation of state, since the stochastic dynamics respects the conservation law $N=1$ (see section II.1 and appendix E).
@@
$H_{S}$ $=$ $\sum_{k=1}^{L}\left(2r_{k}^{2}-r_{k}r_{k-1}-r_{k}r_{k+1}-\frac{\alpha}{2}r_{k}^{4}\right)\rightarrow\mbox{min.},$ (50)
$\sum_{k=1}^{L}r_{k}^{2}=1\,.$
@@
The ground state of the Hamiltonian with energy $E_{\mathrm{low}}(\alpha)$ corresponds to the zero-temperature state, $\beta=\infty$. Since our Hamiltonian is bounded, the state with highest energy $E_{\mathrm{up}}(\alpha)$ corresponds to the high-temperature limit $\beta=$-$\infty$ which means a negative-temperature state where one approaches zero temperature from below. Such features are quite common in systems with bounded energy, e.g., spin systems or Laser systems (population inversion is a negative-temperature state, which is unstable in experiments since atoms are coupled to a positive temperature, i.e., colder heat bath), but they do not occur in conventional mechanical systems, as the kinetic energy is unbounded. In the latter, the highest-temperature state corresponds to $\beta=0$, i.e., the positive temperature tending to infinity. In the setup of the DNSE with conserved normalisation $N=1$ the infinite-temperature state $\beta=0$ is intermediate between the low-temperature state $\beta=\infty$ and the high-temperature state $\beta=$-$\infty$. The corresponding distribution in phase space is uniform on the sphere determined by the conservation law $N=1$, see eq. (9), and the corresponding expectation values can be easily computed (see for instance [EBRAHIMI2025134905]) to be $\langle|c_{k}|^{2}\rangle=1/L$, $\langle|c_{k}|^{4}\rangle=2/(L(L+1)$, and $\langle c_{k}\bar{c}_{k-1}\rangle=0$. The expectation of the Hamiltonian (43) then defines the infinite-temperature line in the phase diagram
@@
Case $\alpha<0$: We can deal with this case either by employing the symmetry discussed in section II.2 or by a simple calculation along the lines of the previous case.
@@
In summary, these considerations can be condensed into a phase diagram which shows the accessible region in the $E$-$\alpha$ parameter plane together with positive and negative temperature domains (see figure 9). The phase diagram displays the symmetry mentioned in section II.2.
Figure 9: Phase diagram of accessible energies $E$ and on-site potential strengths $\alpha$ (shaded regions) for the Hamiltonian (1) with system size $L=64$. Blue: boundaries $E_{up}(\alpha)$ for $\alpha>0$ (see eq. (47)) and $E_{\mathrm{low}}(\alpha)$ for $\alpha<0$ (see eq. (53)) caused by ferromagnetic and antiferromagnetic states. Black: boundaries $E_{\mathrm{low}}(\alpha)$ for $\alpha>0$ caused by positive breather and $E_{\mathrm{up}}(\alpha)$ for $\alpha<0$ caused by a staggered breather. Cyan and red (solid/dashed): analytic upper and lower bounds for the boundary caused by breather states (see eq. (49) for $\alpha>0$ and eq. (55) for $\alpha<0$). Bronze: infinite temperature line (see eq. (52)). Light shading indicates the region of predominantly localised breather-type states, dark shading the region of predominantly antiferromagnetic ($\alpha>0$) or ferromagnetic ($\alpha<0$) states. The different types of shading also correlates with positive and negative temperature regions (as indicated by labels). Yellow: the line $\alpha=0$ for which an explicit analytic expression for the caloric equation of state, $E(\beta)$, is available (see eq. (100)). Green: Phase transition line, eq. (16), obtained by mean field theory (see section IV for details). The green dots are numerical results for the breakdown of the localised phase obtained from simulations in a system of size $L=1024$ with noise strength $\sigma=0.3$, using time series of length $6.8\times 10^{5}$.
@@ Appendix C Fokker-Planck equation and stationary distribution
$J_{c_{k}}(\rho_{t})$ $=$ $\frac{\sigma^{2}}{2}c_{k}^{2}\left(\beta\frac{\partial H_{S}}{\partial c_{k}}\rho_{t}+\frac{\partial\rho_{t}}{\partial c_{k}}\right)$
$-\frac{\sigma^{2}}{2}|c_{k}|^{2}\left(\beta\frac{\partial H_{S}}{\partial\bar{c}_{k}}\rho_{t}+\frac{\partial\rho_{t}}{\partial\bar{c}_{k}}\right)$
@@
$c_{k}^{2}\frac{\partial\rho_{t}f(N)}{\partial c_{k}}-|c_{k}|^{2}\frac{\partial\rho_{t}f(N)}{\partial\bar{c}_{k}}$ (60)
$+\rho_{t}f^{\prime}(N)\left(c_{k}^{2}\bar{c}_{k}-|c_{k}|^{2}c_{k}\right)$
@@
$\bar{c}_{k}J_{c_{k}}(\rho_{t})+c_{k}J_{\bar{c}_{k}}(\rho_{t})$ $=$ $\frac{\sigma^{2}}{2}\left(\beta\frac{\partial H_{S}}{\partial c_{k}}+\frac{\partial\rho_{t}}{\partial c_{k}}\right)\left(\bar{c}_{k}c_{k}^{2}-c_{k}|c_{k}|^{2}\right)+\frac{\sigma^{2}}{2}\left(\beta\frac{\partial H_{S}}{\partial\bar{c}_{k}}+\frac{\partial\rho_{t}}{\partial\bar{c}_{k}}\right)\left(c_{k}\bar{c}_{k}^{2}-\bar{c}_{k}|c_{k}|^{2}\right)=0$ (62)
@@
$\frac{\partial\rho_{t}f(N)}{\partial t}$ $=$ $-f(N)\{H_{S},\rho_{t}\}$
$-f(N)\sum_{k=1}^{L}\left(\frac{\partial J_{c_{k}}(\rho_{t})}{\partial c_{k}}+\frac{\partial J_{\bar{c}_{k}}(\rho_{t})}{\partial\bar{c}_{k}}\right)$
$=$ $-\{H_{S},\rho_{t}f(N)\}$
$-\sum_{k=1}^{L}\left(\frac{\partial J_{c_{k}}(\rho_{t}f(N))}{\partial c_{k}}+\frac{\partial J_{\bar{c}_{k}}(\rho_{t}f(N))}{\partial\bar{c}_{k}}\right)$
@@
Finally, eq. (58) and the stationary solution (65) admit an H-theorem, as easily seen, when adjusting the standard reasoning [Risk:89] to our particular setup. Consider the expression
@@
$L(t)=\int_{\mathbb{C}^{L}}\Big[\frac{e^{-\beta H_{S}}}{\rho_{t}}\ln\!\frac{e^{-\beta H_{S}}}{\rho_{t}}$
$-\frac{e^{-\beta H_{S}}}{\rho_{t}}+1\Big]\rho_{t}\,f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})>0.$ (67)
if $\rho_{t}f(N)\neq\rho_{\beta}$ and $L(t)=0$ if $\rho_{t}f(N)=\rho_{\beta}$, that means $L(t)$ has a global minimum. Furthermore
@@
$=$ $\int_{\mathbb{C}^{L}}e^{-\beta H_{S}}\{H_{S},\ln\rho_{t}\}f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})+\sum_{k=1}^{L}\int_{\mathbb{C}^{L}}\frac{e^{-\beta H_{S}}}{\rho_{t}}\left(\frac{\partial J_{c_{k}}(\rho_{t})}{\partial c_{k}}+\frac{\partial J_{\bar{c}_{k}}(\rho_{t})}{\partial\bar{c}_{k}}\right)f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})$
@@
$J_{c_{k}}(\rho_{t})=\frac{\sigma^{2}}{2}\left(c_{k}^{2}\frac{\partial\rho_{t}/e^{-\beta H_{2}}}{\partial c_{k}}-|c_{k}|^{2}\frac{\partial\rho_{t}/e^{-\beta H_{2}}}{\partial\bar{c}_{k}}\right)$ (69)
@@
$\dot{L}(t)=-\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\int_{\mathbb{C}^{L}}\left(\frac{e^{-\beta H_{S}}}{\rho_{t}}\right)^{2}(\bar{p}_{k},p_{k})\left(\begin{array}[]{rr}1&-1\\ -1&1\end{array}\right)\left(\begin{array}[]{c}p_{k}\\ \bar{p}_{k}\end{array}\right)f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})$ (70)
@@ Appendix D Numerical integration scheme
$\frac{\partial\mathbf{U}(t_{0},t)}{\partial t}=\mathbf{U}(t_{0},t)\left(i\mathbf{L}_{A}+i\mathbf{L}_{B}(t)-\mathbf{\Lambda}\right),$
$\mathbf{U}(t_{0},t_{0})=\mathbf{1}$ (72)
@@
$\mathbf{\Lambda}F$ $=$ $\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}c_{k}\left(c_{k}(\bar{c}_{k+1}+\bar{c}_{k-1})-\bar{c}_{k}(c_{k+1}+c_{k-1})\right)\frac{\partial F}{\partial c_{k}}+\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\bar{c}_{k}\left(\bar{c}_{k}(c_{k+1}+c_{k-1})-c_{k}(\bar{c}_{k+1}+\bar{c}_{k-1})\right)\frac{\partial F}{\partial\bar{c}_{k}}\,.$
@@
$\mathbf{U}(t_{0},t_{0}+\tau)=\exp\left(i\mathbf{L}_{B}^{<}\tau\right)\exp\left(i\mathbf{L}_{A}\tau-\mathbf{\Lambda}\tau\right)$
$\exp\left(i\mathbf{L}_{B}^{>}\tau\right)+\mathcal{O}(\tau^{3})$ (75)
@@
$\langle D_{n}^{<}D_{m}^{<}\rangle=\frac{1}{3}\delta_{n,m},\qquad\langle D_{n}^{>}D_{m}^{>}\rangle=\frac{1}{3}\delta_{n,m},$
$\langle D_{n}^{<}D_{m}^{>}\rangle=\frac{1}{6}\delta_{n,m}\,.$ (77)
Using, for instance, the usual series expansion, we can finally write the combined exponential in eq. (D) as a product, down to terms of second order
$\exp\left(i\mathbf{L}_{A}\tau-\mathbf{\Lambda}\tau\right)=\exp\left(-\mathbf{\Lambda}\tau/2\right)\exp\left(i\mathbf{L}_{A}\tau\right)$
$\exp\left(-\mathbf{\Lambda}\tau/2\right)+\mathcal{O}(\tau^{3})\,.$ (78)
Eqs. (D) and (D) define a numerical integration step with stepsize $\tau$, where each integration step consists of five individual transformations. The action of the exponentials $\exp(i\mathbf{L}_{A}\tau)$ and $\exp(i\mathbf{L}_{B}^{\lessgtr}\tau)$ can be calculated in closed analytic form, as the corresponding dynamics is governed by a system of independent entities or by a nearest neighbour coupled harmonic chain, cf. eqs. (D) and (76) (see [EBRAHIMI2025134905] for the explicit analytic expressions). In addition, those transformations are symplectic and they preserve the normalisation so that these parts meet the constraints mentioned above. We still need to discuss the effect of the dissipative part $\exp(-\mathbf{\Lambda}\tau/2)$.
@@
$\dot{c}_{n}(t)=-\beta\frac{\sigma^{2}}{2}c_{n}(t)\Big[c_{n}(t)\big(\bar{c}_{\,n-1}(t)+\bar{c}_{\,n+1}(t)\big)$
$\quad-\bar{c}_{n}(t)\big(c_{\,n-1}(t)+c_{\,n+1}(t)\big)\Big].$ (79)
By construction, this set of equations just contains the dissipative part of the original SDNSE, eq. (11). That means we are now going to discuss the sole impact of the damping on the motion. It is fairly straightforward to show that eq. (D) preserves the amplitudes $|c_{n}|^{2}$ since
@@
Such a property already implies that the action of $\exp(-\mathbf{\Lambda}\tau/2)$ preserves the normalisation since each individual amplitude $|c_{n}^{2}|$ remains constant. In addition, we can conclude that the time dependent solutions of eq. (D) have a constant absolute value
@@
where the time dependent complex phases obey (see eq. (D))
$\dot{\theta}_{n}(t)=\beta\sigma^{2}|c_{n}|\Big[|c_{n-1}|\sin\!\left(\theta_{n-1}(t)-\theta_{n}(t)\right)$
$\qquad+|c_{n+1}|\sin\!\left(\theta_{n+1}(t)-\theta_{n}(t)\right)\Big].$ (82)
Hence, the damping mechanism on its own only readjusts the complex phases but leaves the individual amplitudes untouched. Eq. (D) has the form of a phase coupled oscillator model, like the Kuramoto model, with nearest neighbour couplings determined by the constant amplitudes. While no analytic closed-form solution is available in these cases, we may compute approximate values for $\theta_{n}(\tau/2)$ by a single numerical integration step of second order accuracy, e.g., by a step of the midpoint method. Even with these numerical errors the resulting values for $c_{n}(\tau/2)=\exp(-\mathbf{\Lambda}\tau/2)c_{n}=|c_{n}|\exp(i\theta_{n}(\tau/2))$ will preserve normalisation, so that our scheme fulfils all the requirements mentioned above.

### v2

@@ Appendix B Phase diagram of the DNSE
The phase diagram of the DNSE, i.e., the bounds for the energy and a representation of the different regimes in the parameter plane, is well established in the literature. The considerations are often based on a standardised version of the DNSE (see, for instance, [rasmussen_statistical_2000, rumpf_simple_2004, BaIuLiVu_PR21]) and those results can be easily translated to our setup using suitable parameter transformations (see e.g. [EBRAHIMI2025134905] for the transformation formulas). For the convenience of the reader, we here summarise the elementary estimates within our notation of eq. (1). The two parameters in our setup are the energy $E=H_{S}$ and the on-site potential strength $\alpha$. While the phase diagram summarises the properties of $H_{S}$ subjected to the constraint $N=1$, the diagram also applies for the SDNSE, eq. (11), if one is able to translate between temperature and energy using the caloric equation of state, since the stochastic dynamics respects the conservation law $N=1$ (see section 2.1 and appendix E).
@@
$H_{S}$ $=$ $\sum_{k=1}^{L}\left(2r_{k}^{2}-r_{k}r_{k-1}-r_{k}r_{k+1}-\frac{\alpha}{2}r_{k}^{4}\right)\rightarrow\mbox{min.},\sum_{k=1}^{L}r_{k}^{2}=1\,.$ (49)
@@
The ground state of the Hamiltonian with energy $E_{\mathrm{low}}(\alpha)$ corresponds to the zero-temperature state, $\beta=\infty$. Since our Hamiltonian is bounded, the state with highest energy $E_{\mathrm{up}}(\alpha)$ corresponds to the high-temperature limit $\beta=$-$\infty$ which means a negative-temperature state where one approaches zero temperature from below. Such features are quite common in systems with bounded energy, e.g., spin systems or Laser systems (population inversion is a negative-temperature state, which is unstable in experiments since atoms are coupled to a positive temperature, i.e., colder heat bath), but they do not occur in conventional mechanical systems, as the kinetic energy is unbounded. In the latter, the highest-temperature state corresponds to $\beta=0$, i.e., the positive temperature tending to infinity. In the setup of the DNSE with conserved normalisation $N=1$ the infinite-temperature state $\beta=0$ is intermediate between the low-temperature state $\beta=\infty$ and the high-temperature state $\beta=$-$\infty$. The corresponding distribution in phase space is uniform on the sphere determined by the conservation law $N=1$, see eq. (9), and the corresponding expectation values can be easily computed (see for instance [EBRAHIMI2025134905]) to be $\langle|c_{k}|^{2}\rangle=1/L$, $\langle|c_{k}|^{4}\rangle=2/(L(L+1))$, and $\langle c_{k}\bar{c}_{k-1}\rangle=0$. The expectation of the Hamiltonian (42) then defines the infinite-temperature line in the phase diagram
@@
Case $\alpha<0$: We can deal with this case either by employing the symmetry discussed in section 2.2 or by a simple calculation along the lines of the previous case.
@@
In summary, these considerations can be condensed into a phase diagram which shows the accessible region in the $E$-$\alpha$ parameter plane together with positive and negative temperature domains (see figure 9). The phase diagram displays the symmetry mentioned in section 2.2.
Figure 9: Phase diagram of accessible energies $E$ and on-site potential strengths $\alpha$ (shaded regions) for the Hamiltonian (1) with system size $L=64$. Blue: boundaries $E_{up}(\alpha)$ for $\alpha>0$ (see eq. (46)) and $E_{\mathrm{low}}(\alpha)$ for $\alpha<0$ (see eq. (52)) caused by ferromagnetic and antiferromagnetic states. Black: boundaries $E_{\mathrm{low}}(\alpha)$ for $\alpha>0$ caused by positive breather and $E_{\mathrm{up}}(\alpha)$ for $\alpha<0$ caused by a staggered breather (this solution was calculated by solving the minimisation problem, eq. (49), by numerical means). Cyan and red (solid/dashed): analytic upper and lower bounds for the boundary caused by breather states (see eq. (48) for $\alpha>0$ and eq. (54) for $\alpha<0$). Bronze: infinite temperature line (see eq. (51)). Light shading indicates the region of predominantly localised breather-type states, dark shading the region of predominantly antiferromagnetic ($\alpha>0$) or ferromagnetic ($\alpha<0$) states. The different types of shading also correlates with positive and negative temperature regions (as indicated by labels). Yellow: the line $\alpha=0$ for which an explicit analytic expression for the caloric equation of state, $E(\beta)$, is available (see eq. (99)). Green: Phase transition line, eq. (16), obtained by mean field theory (see section 4 for details). The green dots are numerical results for the breakdown of the localised phase obtained from simulations in a system of size $L=1024$ with noise strength $\sigma=0.3$, using time series of length $6.8\times 10^{5}$.
@@ Appendix C Fokker-Planck equation and stationary distribution
$J_{c_{k}}(\rho_{t})$ $=$ $\frac{\sigma^{2}}{2}c_{k}^{2}\left(\beta\frac{\partial H_{S}}{\partial c_{k}}\rho_{t}+\frac{\partial\rho_{t}}{\partial c_{k}}\right)-\frac{\sigma^{2}}{2}|c_{k}|^{2}\left(\beta\frac{\partial H_{S}}{\partial\bar{c}_{k}}\rho_{t}+\frac{\partial\rho_{t}}{\partial\bar{c}_{k}}\right)$
@@
$c_{k}^{2}\frac{\partial\rho_{t}f(N)}{\partial c_{k}}-|c_{k}|^{2}\frac{\partial\rho_{t}f(N)}{\partial\bar{c}_{k}}$ $=$ $f(N)\left(c_{k}^{2}\frac{\partial\rho_{t}}{\partial c_{k}}-|c_{k}|^{2}\frac{\partial\rho_{t}}{\partial\bar{c}_{k}}\right)+\rho_{t}f^{\prime}(N)\left(c_{k}^{2}\bar{c}_{k}-|c_{k}|^{2}c_{k}\right)$ (59)
@@
$\bar{c}_{k}J_{c_{k}}(\rho_{t})+c_{k}J_{\bar{c}_{k}}(\rho_{t})$ $=$ $\frac{\sigma^{2}}{2}\left(\beta\frac{\partial H_{S}}{\partial c_{k}}+\frac{\partial\rho_{t}}{\partial c_{k}}\right)\left(\bar{c}_{k}c_{k}^{2}-c_{k}|c_{k}|^{2}\right)$ (61)
$+\frac{\sigma^{2}}{2}\left(\beta\frac{\partial H_{S}}{\partial\bar{c}_{k}}+\frac{\partial\rho_{t}}{\partial\bar{c}_{k}}\right)\left(c_{k}\bar{c}_{k}^{2}-\bar{c}_{k}|c_{k}|^{2}\right)=0$
@@
$\frac{\partial\rho_{t}f(N)}{\partial t}$ $=$ $-f(N)\{H_{S},\rho_{t}\}-f(N)\sum_{k=1}^{L}\left(\frac{\partial J_{c_{k}}(\rho_{t})}{\partial c_{k}}+\frac{\partial J_{\bar{c}_{k}}(\rho_{t})}{\partial\bar{c}_{k}}\right)$ (63)
$=$ $-\{H_{S},\rho_{t}f(N)\}-\sum_{k=1}^{L}\left(\frac{\partial J_{c_{k}}(\rho_{t}f(N))}{\partial c_{k}}+\frac{\partial J_{\bar{c}_{k}}(\rho_{t}f(N))}{\partial\bar{c}_{k}}\right)$
@@
Finally, eq. (57) and the stationary solution (64) admit an H-theorem. For that purpose we adjust the standard reasoning, see for instance [Risk:89], and introduce a relative entropy $L(t)$, sometimes called the Kullback-Leibler divergence, which will serve as a Lyapunov function. Consider the expression
@@
$L(t)=\int_{\mathbb{C}^{L}}\Big[\frac{e^{-\beta H_{S}}}{\rho_{t}}\ln\!\frac{e^{-\beta H_{S}}}{\rho_{t}}-\frac{e^{-\beta H_{S}}}{\rho_{t}}+1\Big]\rho_{t}\,f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})\geq 0.$ (66)
In fact $L(t)>0$ if $\rho_{t}f(N)\neq\rho_{\beta}$ and $L(t)=0$ if $\rho_{t}f(N)=\rho_{\beta}$ so that $L(t)$ has a global minimum at $\rho_{t}f(N)=\rho_{\beta}$. Furthermore
@@
$=$ $\int_{\mathbb{C}^{L}}e^{-\beta H_{S}}\{H_{S},\ln\rho_{t}\}f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})$
$+\sum_{k=1}^{L}\int_{\mathbb{C}^{L}}\frac{e^{-\beta H_{S}}}{\rho_{t}}\left(\frac{\partial J_{c_{k}}(\rho_{t})}{\partial c_{k}}+\frac{\partial J_{\bar{c}_{k}}(\rho_{t})}{\partial\bar{c}_{k}}\right)f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})$
@@
$J_{c_{k}}(\rho_{t})=\frac{\sigma^{2}}{2}e^{-\beta H_{S}}\left(c_{k}^{2}\frac{\partial\rho_{t}/e^{-\beta H_{S}}}{\partial c_{k}}-|c_{k}|^{2}\frac{\partial\rho_{t}/e^{-\beta H_{S}}}{\partial\bar{c}_{k}}\right)$ (68)
@@
$\dot{L}(t)=-\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\int_{\mathbb{C}^{L}}\left(\frac{e^{-\beta H_{S}}}{\rho_{t}}\right)^{2}(\bar{p}_{k},p_{k})\left(\begin{array}[]{rr}1&-1\\ -1&1\end{array}\right)\left(\begin{array}[]{c}p_{k}\\ \bar{p}_{k}\end{array}\right)e^{-\beta H_{S}}f(N)\prod_{k=1}^{L}d(c_{k},\bar{c}_{k})$ (69)
@@ Appendix D Numerical integration scheme
$\frac{\partial\mathbf{U}(t_{0},t)}{\partial t}=\mathbf{U}(t_{0},t)\left(i\mathbf{L}_{A}+i\mathbf{L}_{B}(t)-\mathbf{\Lambda}\right),\qquad\mathbf{U}(t_{0},t_{0})=\mathbf{1}$ (71)
@@
$\mathbf{\Lambda}F$ $=$ $\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}c_{k}\left(c_{k}(\bar{c}_{k+1}+\bar{c}_{k-1})-\bar{c}_{k}(c_{k+1}+c_{k-1})\right)\frac{\partial F}{\partial c_{k}}$
$+\beta\frac{\sigma^{2}}{2}\sum_{k=1}^{L}\bar{c}_{k}\left(\bar{c}_{k}(c_{k+1}+c_{k-1})-c_{k}(\bar{c}_{k+1}+\bar{c}_{k-1})\right)\frac{\partial F}{\partial\bar{c}_{k}}\,.$
@@
$\mathbf{U}(t_{0},t_{0}+\tau)=\exp\left(i\mathbf{L}_{B}^{<}\tau\right)\exp\left(i\mathbf{L}_{A}\tau-\mathbf{\Lambda}\tau\right)\exp\left(i\mathbf{L}_{B}^{>}\tau\right)+\mathcal{O}(\tau^{3})$ (74)
@@
$\langle D_{n}^{<}D_{m}^{<}\rangle=\frac{1}{3}\delta_{n,m},\qquad\langle D_{n}^{>}D_{m}^{>}\rangle=\frac{1}{3}\delta_{n,m},\qquad\langle D_{n}^{<}D_{m}^{>}\rangle=\frac{1}{6}\delta_{n,m}\,.$ (76)
Using, for instance, the usual series expansion, we can finally write the combined exponential in eq. (74) as a product, down to terms of second order
$\exp\left(i\mathbf{L}_{A}\tau-\mathbf{\Lambda}\tau\right)=\exp\left(-\mathbf{\Lambda}\tau/2\right)\exp\left(i\mathbf{L}_{A}\tau\right)\exp\left(-\mathbf{\Lambda}\tau/2\right)+\mathcal{O}(\tau^{3})\,.$ (77)
Eqs. (74) and (77) define a numerical integration step with stepsize $\tau$, where each integration step consists of five individual transformations. The action of the exponentials $\exp(i\mathbf{L}_{A}\tau)$ and $\exp(i\mathbf{L}_{B}^{\lessgtr}\tau)$ can be calculated in closed analytic form, as the corresponding dynamics is governed by a system of independent entities or by a nearest neighbour coupled harmonic chain, cf. eqs. (D) and (75) (see [EBRAHIMI2025134905] for the explicit analytic expressions). In addition, those transformations are symplectic and they preserve the normalisation so that these parts meet the constraints mentioned above. We still need to discuss the effect of the dissipative part $\exp(-\mathbf{\Lambda}\tau/2)$.
@@
$\dot{c}_{n}(t)=-\beta\frac{\sigma^{2}}{2}c_{n}(t)\Big[c_{n}(t)\big(\bar{c}_{\,n-1}(t)+\bar{c}_{\,n+1}(t)\big)-\bar{c}_{n}(t)\big(c_{\,n-1}(t)+c_{\,n+1}(t)\big)\Big].$ (78)
By construction, this set of equations just contains the dissipative part of the original SDNSE, eq. (11). That means we are now going to discuss the sole impact of the damping on the motion. It is fairly straightforward to show that eq. (78) preserves the weights $|c_{n}|^{2}$ since
@@
Such a property already implies that the action of $\exp(-\mathbf{\Lambda}\tau/2)$ preserves the normalisation since each individual weight $|c_{n}^{2}|$ remains constant. In addition, we can conclude that the time dependent solutions of eq. (78) have a constant absolute value
@@
where the time dependent complex phases obey (see eq. (78))
$\dot{\theta}_{n}(t)=\beta\sigma^{2}|c_{n}|\Big[|c_{n-1}|\sin\!\left(\theta_{n-1}(t)-\theta_{n}(t)\right)+|c_{n+1}|\sin\!\left(\theta_{n+1}(t)-\theta_{n}(t)\right)\Big].$ (81)
Hence, the damping mechanism on its own only readjusts the complex phases but leaves the individual $|c_{n}|$ untouched. Eq. (81) has the form of a phase coupled oscillator model, like the Kuramoto model, with nearest neighbour couplings determined by the constant values of the $|c_{n}|$. While no analytic closed-form solution is available in these cases, we may compute approximate values for $\theta_{n}(\tau/2)$ by a single numerical integration step of second order accuracy, e.g., by a step of the midpoint method. Even with these numerical errors the resulting values for $c_{n}(\tau/2)=\exp(-\mathbf{\Lambda}\tau/2)c_{n}=|c_{n}|\exp(i\theta_{n}(\tau/2))$ will preserve normalisation, so that our scheme fulfils all the requirements mentioned above.